Circular No. 7419

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

VARIABLE STAR IN NGC 3432 H. Yamaoka, Kyushu University, reports that digitized scans of R

plates (limiting mag ∼ 20) from the second Palomar Sky Survey taken on 1998 May 16 and 17 show what appears to be a point source at mag ∼ 19.5 embedded in an H II region at a position close to the location of the new variable (cf. IAUC 7415): α = 10^{h}52^{m}41^{s}.29, δ = +36^{o}40′08′′.8 (equinox 2000.0; measurement made with 44 USNO-A2.0 reference stars). The object does not appear on other Digital Sky Survey images (B_{j} plates taken in 1991 and 1996, R plate taken in 1953).

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD S. Lawrence, B. Sugerman, and A. Crotts, Columbia University, report

on the development of new structure and brightening in the inner equatorial ring of SN 1987A: “Hubble Space Telescope STIS F28X50LP imaging and G750M spectroscopy (52′′ × 2′′ slit) were obtained on May 1 and compared to earlier STIS observations. The first hot spot (IAUC 6665) now appears to be moderately resolved, with a FWHM ∼ 15 percent wider than stars in both images and spectra, and a faint, asymmetric extension of the Hα emission appears eastward from the spot. These both require confirmation with high-resolution imaging. All previously reported spots (IAUC 7354, 7359, 7360) are detected in Hα; the first four are also detected in [O I] 630.0nm; and the first spot is detected in [N II] (654.8, 658.3 nm), He I (667.8 nm), and marginally in [S II] (673.1 nm). A new spot is seen in Hα at p.a. 285^{o}, and additional flux is suggested over p.a. 30^{o}–90^{o} but is overwhelmed by the first four spots. The bulk of the emission from most of the spots is Doppler-shifted in the same sense as the equatorial ring, but motion is detected counter to this, particularly in the first spot, with redshifts up to +200 km/s (and blueshifts to −275 km/s, measured at 10 percent of peak line flux). Fluxes measured from the G750M spectrum (once equatorialring flux is subtracted, using a similar, scaled spectrum from 1997 Apr.

27) yield the following values for all seven hot spots, listed in discovery order (fluxes in units of 10^{−15} erg s^{−1} cm^{−2}, followed by position angle for clarity): 14.8, 29^{o}; 3.9, 106^{o}; 2.9, 126^{o}; 0.8, 91^{o}; 0.9, 139^{o}; 1.6, 229^{o}; 0.6, 285^{o}.”

Corrigendum. On IAUC 7359, line 10, for erg s^{−1} cm^{−2}. read erg cm^{−2} s^{−1} arcsec^{−2}.

2000 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green